The surface area of a cylinder with height, h, and base radius, r, is given by
A = 2 rh + 2 r2. Using 3.14 for , find the surface area of a cylinder with radius of 4 cm and a height of 20 cm.

Answer by checkley77(12569)

A = pi*2r*h) + 2(pi*r^2)
radius of 4 cm and a height of 20 cm.
A=3.14*8*20+2*3.14*16
A=502.4+100.48
A=602.88 cm^2 is the total surface area.